OpenStudy (anonymous):

how do you solve \sqrt{x+4}=0

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

i got that part but i jus need the steps

OpenStudy (anonymous):

Since sqrt(x)=0 when x=0

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

(sqrt(x+4)^2)=0^2 x+4=0 x=-4